The electric field due to a charged particle at a point 1.0 m away from it has
magnitude 81NC-1.What is the magnitude of the electric charge on the particle?
What is the magnitude of the electrostatic force on a particle having the same
charge kept at a distance of 2.0 m from it?
1
Expert's answer
2020-03-02T09:56:16-0500

The electric field due to a charged particle according to Coulomb's law is "E=k\frac {q}{r^2}" where "k=9\cdot10^9Nm^2C^{-2}". Thus "q=\frac{E\cdot r^2}{k}=\frac{81 NC^{-1}\cdot1m^2}{9\cdot 10^9 Nm^2C^{-2}}=9\cdot 10^{-9}C" is chage of the particle. The magnitude of the electrostatic force between two particles with chage "q" at "R=2.0m" is

"F=k\frac{q^2}{R^2}=9\cdot 10^9\frac{81\cdot 10^{-18}}{4}=1.8\cdot 10^{-7}N"

Answer: The magnitude of the electric charge on the particle is "9\cdot 10^{-9}C" . The magnitude of the electrostatic force on a particle having the same charge kept at a distance of 2.0 m from it is "1.8\cdot 10^{-7}N" .